# CDsampling: An R Package for Constrained D-Optimal Sampling in Paid Research Studies

### v2

@@ Abstract
In the context of paid research studies and clinical trials, budget considerations often require patient sampling from available populations which comes with inherent constraints. We introduce the R package CDsampling, which is the first to our knowledge to integrate optimal design theories within the framework of constrained sampling. This package offers the possibility to find both D-optimal approximate and exact allocations for samplings with or without constraints. Additionally, it provides functions to find constrained uniform sampling as a robust sampling strategy when the model information is limited. To demonstrate its efficacy, we provide simulated examples and a real-data example with datasets embedded in the package and compare them with classical sampling methods. Furthermore, it revisits the theoretical results of the Fisher information matrix for generalized linear models (including regular linear regression model) and multinomial logistic models, offering functions for its computation.
@@ 1   Introduction
Commonly used sampling strategies include simple random sampling, stratified sampling, and cluster sampling. Simple random sampling is the simplest form of probability sampling, where each element has an equal chance of being selected (Tillé, 2006). Stratified sampling divides the population into homogeneous subgroups and applies random sampling within each subgroup (Lohr, 2019). Cluster sampling splits the population into clusters and randomly selects some clusters as the sample units. However, these methods have their drawbacks. Cluster sampling is relatively low-cost but less precise than simple random sampling. Stratified sampling can produce more efficient estimators of population means, but it requires finding well-defined and relevant subgroups that cover the entire population (Levy and Lemeshow, 2008). Moreover, these existing methods are based on assumptions that may not hold if the model estimation is the primary goal of the paid research study.
In the proposed CDsampling package, we implement the sampling strategy based on optimal design theory, which can improve the model estimation for paid research studies. Optimal design theory is a branch of experimental design that aims to find the best allocation of experimental units to achieve a specific optimality criterion. For example, D-optimality maximizes the determinant of the information matrix, which minimizes the estimators’ expected volume of the joint confidence ellipsoid. A-optimality minimizes the trace of the inverse information matrix, equivalent to minimizing the average of the variances of the estimators. E-optimality minimizes the maximum eigenvalue of the inverse information matrix, which minimizes the largest expected semi-axis of the confidence ellipsoid and protects against the worst possible case (Fedorov, 1972; Atkinson et al., 2007; Yang and Stufken, 2012; Fedorov and Leonov, 2014). In this paper, we focus on D-optimality due to its overall good performance and mathematical simplicity (Zocchi and Atkinson, 1999; Atkinson et al., 2007). To implement the recommended sampling strategy, we develop an R package called CDsampling (namely, Constrained D-optimal sampling), available on CRAN at https://cran.r-project.org/package=CDsampling. To our best knowledge, the CDsampling is the first R package offering the sampling tool with constraints in paid research studies based on optimal design theory. Package sampling implements random samples using different sampling schemes (Tillé and Matei, 2016). The package also provides functions to obtain (generalized) calibration weights, different estimators, as well as some variance estimators. Package SamplingStrata determines the best stratification for a population frame that ensures the minimum sample cost with precision thresholds (Barcaroli, 2014). On the other hand, there are existing R packages for optimal designs. Package AlgDesign finds exact and approximate allocations of optimal designs under D-, A-, and I-criteria (Wheeler and Braun, 2019; Wheeler, 2004). Package OptimalDesign enables users to compute D-, A-, I-, and c-efficient designs with or without replications, restricted design spaces, and under multiple linear constraints (Harman et al., 2016). Package DoE.wrapper enhances designs from different packages (Grömping, 2011b; Groemping et al., 2020), such as response surface designs from package rsm, Latin hypercube samples from packages lhs and DiceDesign, and D-optimal designs from package AlgDesign. Package acebayes find Bayesian optimal experimental designs with approximate coordinate exchange algorithm (Overstall et al., 2017, 2018). Package OPDOE provides functions for optimal designs with polynomial regression models and ANOVA (Grömping, 2011a). These packages provide programming tools for finding samplings or optimal designs under different criteria and models. However, they do not address the specific challenges of practical feasibility in the constrained sampling scheme of paid research studies. The proposed CDsampling package fills this gap by offering an efficient sampling tool to handle general constraints and common parametric models in paid research studies.
@@ 2.1   Constrained lift-one algorithm
Figure 1 provides a concise summary of the lift-one algorithm applied to general parametric models. The detailed algorithm is provided in Algorithm $3$ from the supplementary materials of Huang et al. (2023). We consider a general study or experiment involving covariates $\mathbf{x}_{i}=(x_{i1},\dots,x_{id})^{\top}$, for $i=1,\dots,m$, referred to as experimental settings. In paid research studies, they could be the stratification factors. Here, $m\geq 2$ denotes the number of combinations of covariate levels. Suppose the responses follow a parametric model $M(\mathbf{x}_{i},\boldsymbol{\theta})$, where $\boldsymbol{\theta}\subseteq\mathbb{R}^{p}$ with $p\geq 2$. Under regularity conditions, the Fisher information matrix of the experiment design can be written as $\mathbf{F}=\sum_{i=1}^{m}n_{i}\mathbf{F}_{i}$, where $\mathbf{F}_{i},i=1,\dots,m$ is the Fisher information matrix at $\mathbf{x}_{i}$ and $n_{i}$ is the number of repeated measurements or subjects assigned to the $i$th experimental setting. We usually call $\mathbf{n}=(n_{1},\dots,n_{m})^{\top}$ the exact allocation, where $n=\sum_{i=1}^{m}n_{i}$, and $\mathbf{w}=(w_{1},\dots,w_{m})^{\top}=(n_{1}/n,\dots,n_{m}/n)^{\top}$ the approximate allocation, where $w_{i}\geq 0$ and $\sum_{i=1}^{m}w_{i}=1$ (Kiefer, 1974; Pukelsheim, 2006; Atkinson et al., 2007). The approximate allocation is theoretically more tractable while the exact allocation is practically more useful. The D-optimal design aims to find the optimal allocation that maximizes $f({\mathbf{w}})=|\mathbf{F}({\mathbf{w}})|=|\sum_{i=1}^{m}w_{i}\mathbf{F}_{i}|$. The lift-one algorithm simplifies a complex multivariate optimization problem into a sequence of simpler univariate optimization problems. This is achieved by “lifting” and optimizing one weight $w_{i}$, within the approximate allocation vector $\mathbf{w}$. Specifically, in step $3^{\circ}$ of the algorithm depicted in Figure 1, the determinant of the Fisher information matrix function concerning the lift-one variable is expressed as $f(z)=f(\mathbf{w}_{i}(z))$ where the variable $z$ substitutes for $w_{i}$ in allocation $\mathbf{w}$, and the remaining weights are adjusted proportionally. The updated weight vector is given by
$\mathbf{w}_{i}(z)=\left(\frac{1-z}{1-w_{i}}w_{1},\ldots,\frac{1-z}{1-w_{i}}w_{i-1},z,\frac{1-z}{1-w_{i}}w_{i+1},\ldots,\frac{1-z}{1-w_{i}}w_{m}\right)^{\top}$
@@
In the context of paid research studies, budgetary constraints often necessitate the selection of a subset of participants. We consider the sampling of $n$ subjects from a larger population of $N$, where $n<N$. A typical restriction in such studies is $n_{i}\leq N_{i}$, where $N_{i}$ represents the number of available subjects from the $i$th experimental setting group. This effectively places an upper bound on the sample size for each subgroup or stratum. Additional constraints may include relationships such as $n_{1}+n_{2}+n_{3}+n_{4}\leq 392$ (see the MLM example in Section 3.1), $4n_{1}\geq n_{3}$ (see Example S1.2 in Section S1), etc. The constrained lift-one algorithm seeks the approximate allocation ${\mathbf{w}}$ that maximizes $|{\mathbf{F}}({\mathbf{w}})|$, on a collection of feasible approximate allocations $S\subseteq S_{0}$ where
$S_{0}:=\{(w_{1},\ldots,w_{m})^{\top}\in\mathbb{R}^{m}\mid w_{i}\geq 0,i=1,\ldots,m;\sum_{i=1}^{m}w_{i}=1\}$
The set $S$ is presumed to be either a closed convex set or a finite union of closed convex sets. The framework of the constrained lift-one algorithm is provided in Figure 2. The details of the algorithm can be found in Algorithm 1 of Huang et al. (2023). In the constrained lift-one algorithm, the search for the optimal lift-one weight $z$ in step $3^{\circ}$ of Figure 2 is confined within the interval of $[r_{i1},r_{i2}]$. To ensure that the resulting allocation is D-optimal, two additional decision steps, labeled steps $7^{\circ}$ and $8^{\circ}$, are incorporated into the algorithm. The reported allocation in step $10^{\circ}$ is the constrained D-optimal approximate allocation for the study. To illustrate the difference between the lift-one algorithm and the constrained lift-one algorithm, we provide examples in Supplementary Section S1.
@@ 2.2   Fisher information for generalized linear models
where $g$ is a link function, $\boldsymbol{\beta}=(\beta_{1},\beta_{2},\dots,\beta_{p})^{\top}$ is the parameter vector of interest, and $\mu_{i}$ is the conditional expectation of response $Y_{i}$ given predictors $\mathbf{X}_{i}=\mathbf{h}(\mathbf{x}_{i})=(h_{1}(\mathbf{x}_{i}),\dots,h_{p}(\mathbf{x}_{i}))^{\top}$, where $i=1,\dots,n$ with known and deterministic predictor functions $\mathbf{h}=(h_{1},\dots,h_{p})^{\top}$. There are various link functions could be used, for example, logit link $\eta_{i}=\log\frac{\mu_{i}}{1-\mu_{i}}$; probit link $\eta_{i}=\Phi^{-1}(\mu_{i})$, where $\Phi(\cdot)$ is the normal cumulative distribution function; and complementary log-log link $\eta_{i}=\log\{-\log(1-\mu_{i})\}$. Regular linear models can be considered as GLMs with the identity link function and normal responses. Suppose we have a design with $m$ design points $\mathbf{x}_{1},\dots,\mathbf{x}_{m}$ that has an exact allocation $(n_{1},\dots,n_{m})$ and $\sum_{i}n_{i}=n$. In the statistical literature, an approximate allocation is more common with $(w_{1},\dots,w_{m})=(\frac{n_{1}}{n},\dots,\frac{n_{m}}{n})$ (Kiefer, 1974). The Fisher information matrix $\mathbf{F}$ under GLMs can be written as (McCullagh and Nelder, 1989; Khuri et al., 2006; Stufken and Yang, 2012; Yang et al., 2016):
@@
where $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{m})^{\top}$ is the $m\times p$ design matrix with $\mathbf{X}_{i}=\mathbf{h}(\mathbf{x}_{i})$, $,\mathbf{W}=\text{diag}\{w_{1}\nu_{1},$$\dots,w_{m}\nu_{m}\}$ is a diagonal matrix with $\nu_{i}=\frac{1}{\text{Var}(Y_{i})}(\frac{\partial\mu_{i}}{\partial\eta_{i}})^{2}$. Here $\nu_{i}$ represents how much information the design point $\mathbf{x}_{i}$ contains. The explicit formats of $\nu_{i}$ with different response distributions and link functions can be found in Table 5 of the Supplementary Material of Huang et al. (2023). To calculate the Fisher information matrix $\mathbf{F}$ given the approximate allocation $\mathbf{w}$, we may use the F_func_GLM() function in the CDsampling package. Additionally, the W_func_GLM() function can be used to find the diagonal elements of the matrix $\mathbf{W}$ in the Fisher information matrix (1) of GLM. An example of finding the Fisher information matrix for GLM is provided in the Supplementary of Example S2.1.
@@ 2.3   Fisher information for multinomial logistic model
We assume that the responses $\mathbf{Y}_{i}=(Y_{i1},\dots,Y_{iJ})$ follow a multinomial distribution with probabilities $(\pi_{i1},\dots,\pi_{iJ})$, where $\pi_{ij}=P(Y_{i}=j\mid{\mathbf{x}}_{i})$, $Y_{i}\in\{1,\ldots,J\}$ is the $i$th original categorical response, $j=1,\dots,J$, and $\sum_{j}\pi_{ij}=1$. If the response variables are nominal, in other words, there is no natural ordering among different levels, a commonly used MLM is the baseline-category logit model with the $J$th level selected as the baseline category. If the response variable is ordinal, that is, we have a natural ordering of response levels, there are three commonly used MLMs in the literature: cumulative logit models, adjacent logit models, and continuation-ratio logit models (Bu et al., 2020; Dousti Mousavi et al., 2023; Wang and Yang, 2024).
In addition to different logit models, the proportional odds (po) assumption is an important concept in multinomial logistic models. The po assumption is a parsimonious model assumption, where a model simultaneously uses all $J-1$ logits in a single model with the same coefficients. This means the covariate effect is constant on the odds ratio among different response levels. When the assumption doesn’t hold, the model is referred to as a non-proportional odds (npo) model and has more parameters in it. When the assumption only holds for part of the parameters, the model is referred to as a partial proportional odds (ppo) model. Commonly used multinomial logit models with po, npo, or ppo assumptions can be summarized in a unified matrix form (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020):
@@
where $\mathbf{F}_{i}=(\frac{\partial{\boldsymbol{\pi}_{i}}}{\partial{\boldsymbol{\theta}^{\top}}})^{\top}\text{diag}(\boldsymbol{\pi}_{i})^{-1}\frac{\partial\boldsymbol{\pi}_{i}}{\partial\boldsymbol{\theta}^{\top}}$ with $\frac{\partial\boldsymbol{\pi}_{i}}{\partial\boldsymbol{\theta}^{\top}}=(\mathbf{C}^{\top}\mathbf{D}_{i}^{-1}\mathbf{L})^{-1}$ and $\mathbf{D}_{i}=\text{diag}(\mathbf{L}\boldsymbol{\pi}_{i})$. Explicit forms of $(\mathbf{C}^{\top}\mathbf{D}_{i}^{-1}\mathbf{L})^{-1}$ can be found in Section S.3 of the Supplementary Material of Bu et al. (2020). To calculate the Fisher information matrix $\mathbf{F}$ for the MLM, one may use the function F_func_MLM() in the CDsampling package. An example of finding the Fisher information matrix for MLM is provided in the Supplementary of Example S2.2.

### v3

@@ Abstract
In the context of paid research studies and clinical trials, budget considerations often require patient sampling from available populations which comes with inherent constraints. We introduce the R package CDsampling, which is the first to our knowledge to integrate optimal design theories within the framework of constrained sampling. This package offers the possibility to find both D-optimal approximate and exact allocations for samplings with or without constraints. Additionally, it provides functions to find constrained uniform sampling as a robust sampling strategy when the model information is limited. To demonstrate its efficacy, we provide simulated examples and a real-data example with datasets embedded in the package and compare them with classical sampling methods. Furthermore, the CDsampling package revisits the theoretical results of the Fisher information matrix for generalized linear models (including regular linear regression model) and multinomial logistic models, offering functions for its computation.
@@ 1   Introduction
Consider a constrained sampling problem commonly encountered in paid research studies. Suppose, in a research study to evaluate a new intervention, $N=500$ volunteers register to participate. Upon registration, the investigators collect basic demographic information, such as gender (female or male) and age groups ($18\sim 25$, $26\sim 64$, $65$ and above) for each volunteer. Treating gender and age as stratification factors, we obtain $m=6$ subgroups. However, due to budget limitations, the study could only accommodate $n=200$ participants. Let $N_{i}$ denote the frequency of volunteers within the $i$th subgroup, where $i=1,\dots,m$. We call the integer number of participants sampled from each subgroup, $n_{i}$, as the exact allocation, and the corresponding proportion $n_{i}/N_{i}$ as the approximate allocation. The goal is to select a sample of $200$ participants $\mathbf{n}=(n_{1},\dots,n_{m})$, such that $\sum_{i=1}^{m}n_{i}=200$ from the pool of $N=500$ volunteers to evaluate the intervention effect most accurately, subject to the constraint that no subgroup is oversampled beyond the number of available volunteers within that subgroup, that is, $n_{i}\leq N_{i}$. This constraint is the most commonly encountered in paid research studies (see Section 3.1 for details, and for constraints of other forms, please refer to Section 3.2).
@@
Commonly used sampling strategies include simple random sampling, stratified sampling, and cluster sampling. Simple random sampling is the most straightforward form of probability sampling, where each element has an equal chance of being selected (Tillé, 2006). Proportionally stratified sampling involves dividing the population into homogeneous subgroups, such as gender and age groups, and applying random sampling to sample proportionally within each subgroup (Lohr, 2019). Cluster sampling, on the other hand, splits the population into heterogeneous clusters, for example, based on the locations of volunteers, and randomly selects some clusters as the sample units. However, these methods have their drawbacks. Cluster sampling is relatively low-cost but less precise than simple random sampling. Stratified sampling can produce more efficient estimators of population means, but it requires finding well-defined and relevant subgroups that cover the entire population (Levy and Lemeshow, 2008). Moreover, these existing methods are based on assumptions that may not hold if the model estimation is the primary goal of the paid research study.
[…]
@@ 2.1   Constrained lift-one algorithm
Figure 1 provides a concise summary of the lift-one algorithm applied to general parametric models. The detailed algorithm is provided in Algorithm $3$ from the Supplementary Material of Huang et al. (2023). We consider a general study or experiment involving covariates $\mathbf{x}_{i}=(x_{i1},\dots,x_{id})^{\top}$, for $i=1,\dots,m$, referred to as experimental settings. In paid research studies, these covariates could be the stratification factors such as the gender and age groups in our motivating example. Here, $m\geq 2$ denotes the number of experimental settings, which corresponds to $m=6$, the number of gender and age groups in the motivating example. Suppose the responses follow a parametric model $M(\mathbf{x}_{i},\boldsymbol{\theta})$, where $\boldsymbol{\theta}\subseteq\mathbb{R}^{p}$ with $p\geq 2$. Under regularity conditions, the Fisher information matrix of the experiment design can be written as $\mathbf{F}=\sum_{i=1}^{m}n_{i}\mathbf{F}_{i}$, where $\mathbf{F}_{i},i=1,\dots,m$ is the Fisher information matrix at $\mathbf{x}_{i}$ and $n_{i}$ is the number of subjects allocated to the $i$th experimental setting. In the setting of paid research studies, $n_{i}$ corresponds to the number of subjects sampled from the $i$th subgroup. We usually call $\mathbf{n}=(n_{1},\dots,n_{m})^{\top}$ the exact allocation, where $n=\sum_{i=1}^{m}n_{i}$, and $\mathbf{w}=(w_{1},\dots,w_{m})^{\top}=(n_{1}/n,\dots,n_{m}/n)^{\top}$ the approximate allocation, where $w_{i}\geq 0$ and $\sum_{i=1}^{m}w_{i}=1$ (Kiefer, 1974; Pukelsheim, 2006; Atkinson et al., 2007). The approximate allocation is theoretically more tractable while the exact allocation is practically more useful. In the statistical literature, approximate allocations are more commonly discussed (Kiefer, 1974). The D-optimal design aims to find the optimal allocation that maximizes $f({\mathbf{w}})=|\mathbf{F}({\mathbf{w}})|=|\sum_{i=1}^{m}w_{i}\mathbf{F}_{i}|$. The lift-one algorithm simplifies a complex multivariate optimization problem into a sequence of simpler univariate optimization problems. This is achieved by “lifting” and optimizing one weight $w_{i}$, within the approximate allocation vector $\mathbf{w}$. Specifically, in step $3^{\circ}$ of the algorithm depicted in Figure 1, the determinant of the Fisher information matrix function concerning the lift-one variable is expressed as $f(z)=f(\mathbf{w}_{i}(z))$ where the variable $z$ substitutes for $w_{i}$ in allocation $\mathbf{w}$, and the remaining weights are adjusted proportionally, denoted as $\mathbf{w}_{i}$. The updated weight vector is given by
$\mathbf{w}_{i}(z)=\left(\frac{1-z}{1-w_{i}}w_{1},\ldots,\frac{1-z}{1-w_{i}}w_{i-1},z,\frac{1-z}{1-w_{i}}w_{i+1},\ldots,\frac{1-z}{1-w_{i}}w_{m}\right)^{\top}.$
@@
In the context of paid research studies, budgetary limitations often necessitate the selection of a subset of participants. We consider the sampling of $n$ subjects from a larger population of $N$, where $n<N$. A typical constraint in such studies is $n_{i}\leq N_{i}$, where $N_{i}$ represents the number of available subjects from the $i$th experimental setting group. This effectively places an upper bound on the sample size for each subgroup (or stratum) ensuring the sampling process does not overdraw from any subgroup. Additional constraints may include $n_{1}+n_{2}+n_{3}+n_{4}\leq 392$ (see the MLM example in Section 3.2), $4n_{1}\geq n_{3}$ (see Example S2.2 in Section S2), etc. The constrained lift-one algorithm seeks the approximate allocation ${\mathbf{w}}$ that maximizes $|{\mathbf{F}}({\mathbf{w}})|$, on a collection of feasible approximate allocations $S\subseteq S_{0}$ where
$S_{0}:=\{(w_{1},\ldots,w_{m})^{\top}\in\mathbb{R}^{m}\mid w_{i}\geq 0,i=1,\ldots,m;\sum_{i=1}^{m}w_{i}=1\}.$
The set $S$ is presumed to be either a closed convex set or a finite union of closed convex sets. The framework of the constrained lift-one algorithm is provided in Figure 2. The details of the algorithm can be found in Algorithm 1 of Huang et al. (2023). In the constrained lift-one algorithm, the search for the optimal lift-one weight $z$ in step $3^{\circ}$ of Figure 2 is confined within the interval of $[r_{i1},r_{i2}]$ (see Example S2.2 in Supplementary Material and Section S.3 in Huang et al. (2023) for details of finding $[r_{i1},r_{i2}]$). To ensure that the resulting allocation is D-optimal, two additional decision steps, labeled steps $7^{\circ}$ and $8^{\circ}$, are incorporated into the algorithm. The reported allocation in step $10^{\circ}$ is the constrained D-optimal approximate allocation for the study. To illustrate the difference between the lift-one algorithm and the constrained lift-one algorithm, we provide examples in Supplementary Material Section S2.
@@ 2.2   Fisher information for generalized linear models
where $g$ is a link function, $\boldsymbol{\beta}=(\beta_{1},\beta_{2},\dots,\beta_{p})^{\top}$ is the parameter vector of interest, and $\mu_{i}$ is the conditional expectation of response $Y_{i}$ given predictors $\mathbf{X}_{i}=\mathbf{h}(\mathbf{x}_{i})=(h_{1}(\mathbf{x}_{i}),\dots,h_{p}(\mathbf{x}_{i}))^{\top}$, where $i=1,\dots,n$ with known and deterministic predictor functions $\mathbf{h}=(h_{1},\dots,h_{p})^{\top}$. There are various link functions could be used, for example, logit link $\eta_{i}=\log\frac{\mu_{i}}{1-\mu_{i}}$; probit link $\eta_{i}=\Phi^{-1}(\mu_{i})$, where $\Phi(\cdot)$ is the normal cumulative distribution function; and complementary log-log link $\eta_{i}=\log\{-\log(1-\mu_{i})\}$. Regular linear models can be considered as GLMs with the identity link function and normal responses. Suppose we have a design with $m$ design points $\mathbf{x}_{1},\dots,\mathbf{x}_{m}$ that has an exact allocation $(n_{1},\dots,n_{m})$ where $\sum_{i}n_{i}=n$ and corresponding approximate allocation $(w_{1},\dots,w_{m})=(\frac{n_{1}}{n},\dots,\frac{n_{m}}{n})$. The Fisher information matrix $\mathbf{F}$ under GLMs can be written as (McCullagh and Nelder, 1989; Khuri et al., 2006; Stufken and Yang, 2012; Yang et al., 2016):
@@
where $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{m})^{\top}$ is the $m\times p$ design matrix with $\mathbf{X}_{i}=\mathbf{h}(\mathbf{x}_{i})$, $,\mathbf{W}=\text{diag}\{w_{1}\nu_{1},$$\dots,w_{m}\nu_{m}\}$ is a diagonal matrix with $\nu_{i}=\frac{1}{\text{Var}(Y_{i})}(\frac{\partial\mu_{i}}{\partial\eta_{i}})^{2}$. Here $\nu_{i}$ represents how much information the design point $\mathbf{x}_{i}$ contains. The explicit formats of $\nu_{i}$ with different response distributions and link functions can be found in Table 5 of the Supplementary Material of Huang et al. (2023). To calculate the Fisher information matrix $\mathbf{F}$ given the approximate allocation $\mathbf{w}$, we may use the F_func_GLM() function in the CDsampling package. Additionally, the W_func_GLM() function can be used to find the diagonal elements of the matrix $\mathbf{W}$ in the Fisher information matrix (1) of GLM. An example of finding the Fisher information matrix for GLM is provided in Example S1.1 of Supplementary Material.
@@ 2.3   Fisher information for multinomial logistic model
We assume that the responses $\mathbf{Y}_{i}=(Y_{i1},\dots,Y_{iJ})$ follow a multinomial distribution with probabilities $(\pi_{i1},\dots,\pi_{iJ})$, where $\pi_{ij}=P(Y_{i}=j\mid{\mathbf{x}}_{i})$, $Y_{i}\in\{1,\ldots,J\}$ is the $i$th original categorical response, $j=1,\dots,J$, and $\sum_{j}\pi_{ij}=1$. If the response variables are nominal, in other words, there is no natural ordering among different levels, a commonly used MLM is the baseline-category logit model with the $J$th level selected as the baseline category. If the response variable is ordinal, that is, we have a natural ordering of response levels, there are three commonly used MLMs in the literature: cumulative logit model, adjacent logit model, and continuation-ratio logit model (Bu et al., 2020; Dousti Mousavi et al., 2023; Wang and Yang, 2024).
In addition to different logit models, the proportional odds (po) assumption is an important concept in MLMs. The po assumption is a parsimonious model assumption, where a model simultaneously uses all $J-1$ logits in a single model with the same coefficients. This means the covariate effect is constant on the odds ratio among different response levels. When the assumption doesn’t hold, the model is referred to as a non-proportional odds (npo) model and has more parameters in it. When the assumption only holds for part of the parameters, the model is referred to as a partial proportional odds (ppo) model. Commonly used multinomial logit models with po, npo, or ppo assumptions can be summarized in a unified matrix form (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020):
@@
where $\mathbf{F}_{i}=(\frac{\partial{\boldsymbol{\pi}_{i}}}{\partial{\boldsymbol{\theta}^{\top}}})^{\top}\text{diag}(\boldsymbol{\pi}_{i})^{-1}\frac{\partial\boldsymbol{\pi}_{i}}{\partial\boldsymbol{\theta}^{\top}}$ with $\frac{\partial\boldsymbol{\pi}_{i}}{\partial\boldsymbol{\theta}^{\top}}=(\mathbf{C}^{\top}\mathbf{D}_{i}^{-1}\mathbf{L})^{-1}$ and $\mathbf{D}_{i}=\text{diag}(\mathbf{L}\boldsymbol{\pi}_{i})$. Explicit forms of $(\mathbf{C}^{\top}\mathbf{D}_{i}^{-1}\mathbf{L})^{-1}$ can be found in Section S.3 of the Supplementary Material of Bu et al. (2020). To calculate the Fisher information matrix $\mathbf{F}$ for the MLM, one may use the function F_func_MLM() in the CDsampling package. An example of finding the Fisher information matrix for MLM is provided in Example S1.2 of the Supplementary Material.